# Breaking the Curse of Dimensionality with Distributed Neural Computation

### v1

@@ Abstract
We present a theoretical approach to overcome the curse of dimensionality using a neural computation algorithm which can be distributed across several machines. Our modular distributed deep learning paradigm, termed neural pathways, can achieve arbitrary accuracy while only loading a small number of parameters into GPU VRAM. Formally, we prove that for every error level $\varepsilon>0$ and every Lipschitz function $f:[0,1]^{n}\to\mathbb{R}$, one can construct a neural pathways model which uniformly approximates $f$ to $\varepsilon$ accuracy over $[0,1]^{n}$ while only requiring networks of $\mathcal{O}(\varepsilon^{-1})$ parameters to be loaded in memory and $\mathcal{O}(\varepsilon^{-1}\log(\varepsilon^{-1}))$ to be loaded during the forward pass. This improves the optimal bounds for traditional non-distributed deep learning models, namely ReLU MLPs, which need $\mathcal{O}(\varepsilon^{-n/2})$ parameters to achieve the same accuracy. The only other available deep learning model that breaks the curse of dimensionality is MLPs with super-expressive activation functions. However, we demonstrate that these models have an infinite VC dimension, even with bounded depth and width restrictions, unlike the neural pathways model. This implies that only the latter generalizes. Our analysis is validated experimentally in both regression and classification tasks, demonstrating that our model exhibits superior performance compared to larger centralized benchmarks.
Distributed Computing, Approximation Theory, Deep Learning, VC Dimension.
@@ 1 Introduction
The curse of dimensionality is a core obstruction in the foundations of machine learning. From the theoretical lens, given by the universal approximation theorem, the curse of dimensionality implies that every standard deep learning model approximately implementing an arbitrary almost-everywhere differentiable function on the $n$-dimensional cube $[0,1]^{n}$, up to an error of $\varepsilon>0$, requires at least $\Omega(\varepsilon^{-n/2})$ parameters (Yarotsky, 2018; Kratsios & Papon, 2022; Shen et al., 2021, 2022b). In practice, model upscaling is limited by the available device memory; Video RAM (VRAM) for GPUs. This happens either in training, during the backward pass when computing the Jacobian of the deep learning model, or even in the forward pass if the model is large enough. Currently, the curse of dimensionality is only known to be breakable in either one of two restrictive scenarios:
Table 1: Overview of Main Results: Depth and width of the MLP architecture $\hat{f}_{\theta}$ and the number of active parameters in GPU VRAM (Params. in GPU) required to approximate an $\alpha$-Hölder function. The maximum number of parameters required during a forward pass, for any given input, are also recorded (Forward Pass). The optimal approximation rates for the standard ReLU MLP are derived in Yarotsky (2017); the optimality of which is expressed in terms of VC dimension in Shen et al. (2022b). The rates for the non-Lipschitz continuous “unstable MLP model” are derived in Shen et al. (2022a). The VC dimension of the class of ReLU MLPs with bounded depth and width is computed in Bartlett et al. (2019), the (finite) VC dimension of our neural pathways model, with bounded depth, width, and number of parts, is shown in Proposition 3.6, and the infinite VC dimension of the class of unstable MLPs is shown in Proposition 3.7.
Model Depth Width Params. in (V)RAM Forward Pass VC Dim Ours $\mathcal{O}(\varepsilon^{-1})$ $\mathcal{O}(1)$ ${\color[rgb]{0.09,0.45,0.27}\definecolor[named]{pgfstrokecolor}{rgb}{0.09,0.45,0.27}\pgfsys@color@rgb@stroke{0.09}{0.45}{0.27}\pgfsys@color@rgb@fill{0.09}{0.45}{0.27}{\mathcal{O}(\varepsilon^{-1})}}$ $\mathcal{O}\big{(}\frac{n}{\alpha}\log(1/\varepsilon)/\varepsilon\big{)}$ Finite ReLU MLP ${\color[rgb]{0.82,0.1,0.26}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.1,0.26}\pgfsys@color@rgb@stroke{0.82}{0.1}{0.26}\pgfsys@color@rgb@fill{0.82}{0.1}{0.26}{\mathcal{O}(\varepsilon^{-n/(2\alpha)})}}$ $\mathcal{O}(1)$ ${\color[rgb]{0.82,0.1,0.26}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.1,0.26}\pgfsys@color@rgb@stroke{0.82}{0.1}{0.26}\pgfsys@color@rgb@fill{0.82}{0.1}{0.26}{\mathcal{O}\big{(}\varepsilon^{-\frac{n}{2\alpha}}\big{)}}}$ $\mathcal{O}\big{(}\varepsilon^{-\frac{n}{2\alpha}}\big{)}$ Finite Unstable MLP $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Infinite
Exceptionally Regular Target Functions. The curse of dimensionality can only be broken when approximating members of highly restrictive classes of target functions (Barron, 1993; Mhaskar et al., 2017; Suzuki, 2019; Opschoor et al., 2022; Galimberti et al., 2022; Gonon et al., 2023). However, there are generally no verifiable guarantees that a target function encountered in practice has the necessary structure for these desired approximation theorems to hold.
@@
Generalization Problems for Super-expressive Multilayer Perceptions (MLPs). Deep learning models with highly oscillatory “super-expressive” activation functions (Yarotsky & Zhevnerchuk, 2020; Yarotsky, 2021; Zhang et al., 2022) boast approximation rates which are not cursed by dimensionality. However, as we will see in Proposition 3.7, many of these models have an infinite VC dimension even when they are restricted to having a bounded depth and width. Their unbounded VC dimension implies that the classifiers implemented by these models, by post-composition with $I_{(0,\infty)}$, do not generalize due to the classical results of Blumer et al. (1989); Hanneke (2016); Brukhim et al. (2022). Thus, the real performance of these models need not achieve the approximation-theoretic optima since they can learn from a finite number of training instances.
This paper demonstrates that distributing neural computation offers a feasible third option for overcoming the curse of dimensionality in deep learning. Importantly, unlike the aforementioned cases, we show that distributed deep learning does not require high regularity of the target function being approximated, nor does it rely on models with unorthodox activation functions. Our deep learning algorithm, termed neural pathways, is distributed across the storage space (i.e., offline memory) of the machine. It only requires a small portion of the model to be loaded into VRAM at any given time, while the entire model is distributed over the machine’s (pragmatically infinite) storage space.
In essence, each point in the input space is assigned to a trainable prototype (or landmark or representative point). Each of these prototypes is linked with a small neural network, specifically designed to approximate the target function for inputs in proximity to its associated prototype. During the forward pass, the input is initially directed to the appropriate prototype, after which the corresponding network is loaded into memory and employed to generate a prediction. Our algorithm follows the principle of route-first and predict-second via locally specialized neural networks. This approach mimics how stimuli activate only specialized neural pathways in the brain, without necessitating the involvement of the entire brain to process any single input.
2 Neural Pathways
In this section, we present the motivation behind the neural pathways framework proposed in this paper, and we describe it from an algorithmic perspective. The primary goal is to provide the reader with an intuitive understanding of the overall procedure before delving into a more involved mathematical analysis in Section 3, which is approached through the lens of approximation theory.
Motivation. Our model formulation is loosely inspired by the cognitive sciences, and more specifically, by the fact that neuroscientists have observed that the brain does not activate completely to process each input, but rather uses specific regions of the overall network at a time, also known as brain functional specialization (Mahon & Cantlon, 2011).
On the other hand, from a computational perspective, our motivation is to increase the size of deep learning models regardless of GPU memory capacity. We theoretically study the effect and feasibility of scaling up models by combining smaller distributed networks specialized to approximate specific regions of the input space. This differs from simply increasing the parameter count of a single large network, which is constrained by available VRAM.
2.1 Model overview
We can visualize our framework as a tree, with each leaf corresponding to a prototype and an associated neural network. In our case, we work with MLPs, but the network need not be a specific architecture. We refer to the combination of each prototype and its associated neural network as a neural pathway within our model. We refer to the overall algorithmic construction as neural pathways in plural. We can think of neural pathways as a framework for neural computation rather than a specific model since this approach could be applied to any other common neural networks such as CNNs (LeCun et al., 1989) or Transformers (Vaswani et al., 2017), for instance. Similarly, we propose specific algorithms to discover the prototypes for our neural pathways, but these could be substituted with enhanced approaches in the future. However, the main goal of this work is not to adapt the framework to a variety of specialized machine learning tasks but rather to provide its initial formulation and, particularly, to analyze its mathematical feasibility in terms of the curse of dimensionality.
The neural pathways algorithm is divided into two stages: first, we discover the prototype locations, and then, we optimize the neural pathways associated with each prototype. In step one, our framework learns how to subdivide the input space, and in step two, we focus on neural network parameter optimization while keeping the input space partitioning fixed. The mathematical soundness of this approach is later presented as part of our main theoretical results in Section 3. A schematic of the algorithm workflow is depicted in Figure 1.
Figure 1: $1$) The distance from each input $x$ to all prototypes $p_{1},\dots,p_{8}$ is queried. $2$) The network ($\hat{f}_{2}$ in the figure) assigned to the nearest prototype ($p_{2}$ in the figure) is loaded onto the GPU and used for prediction.
2.2 Training Algorithm
We now provide an explanation for the training algorithm. As discussed, we break down the algorithm into two parts: discovering the prototypes and training the networks. Conceptually, the prototypes define where in the input space the networks are located, or in other words, where in the input space we expect each of the networks to have the best performance. During inference, we will route a given input to the appropriate network based on its nearest prototype. In essence, each network learns to approximate a specific region of the overall input domain.
Discovering prototypes. In principle, we may not know how to partition the input space. One approach is to utilize standard clustering algorithms like $K$-means, but this might be suboptimal for the downstream task unless we are already operating in a structured latent space, such as those found in pre-trained models (further discussion on this is available in Section 5). Another way is to learn it via gradient descent by optimizing the location of the prototypes for a specific task by following the gradient of the downstream loss. At the beginning of training, we have $\hat{F}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\hat{f}_{1}(x),\dots,\hat{f}_{K}(x))$ which contains a collection of $K$ randomly initialized shallow or small networks (i.e., much smaller than our neural pathways described later). In this first step, we assume that we are able to load all randomly initialized networks into our GPU memory. In particular, this is true because we use small networks with few parameters, which we will later “deepen” in the next step by adding additional hidden layers. We initialize the prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$ randomly from a uniform distribution within the bounds of our training dataset input samples. We use the following expression to train the location of our prototypes $\{p_{k}\}_{k=1}^{K}$ in $\mathbb{R}^{n}$ by minimizing the energy:
where the loss $\ell$ is task-specific; for example, one could use mean squared error for regression and cross-entropy for classification. The $\operatorname{softmax}$ weights the importance of the prediction of each of the neural pathways in $\hat{F}$ for a given input $x$, as a function of the input’s distance to the prototypes, $\|x-p_{i}\|_{i=1}^{K}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\|x-p_{1}\|,...,\|x-p_{K}\|)$. Both the locations of the prototypes and the shallow randomly initialized neural networks assigned to them are optimized.
Deepening the Networks. After the initial training phase, we enhance the networks by incorporating additional layers. Specifically, we introduce linear layers with weights initialized to the identity matrix and bias set to zero, just before the final output layer of each network. To encourage gradient flow in these new layers during the subsequent training stage, we slightly perturb this initialization with small Gaussian noise. This approach is driven by the fact that in the second training stage, each neural pathway can be optimized in a distributed manner. Consequently, we can work with larger networks without the need to load all of them simultaneously into our GPU, allowing for more model parameters. During the first stage, we have already optimized our networks alongside the prototype locations, converging towards a minimum. By initializing the networks with the new layers close to the identity, we can ensure that their output at the start of the second stage of training is similar to that produced by the original networks. This allows us to smoothly continue the optimization process from the point where we previously halted.
Neural Pathway Training. Once prototype locations have been fixed we can independently train neural pathways $\hat{f}_{1},\dots,\hat{f}_{K}$ by minimizing for all $k\in\{1,\dots,K\}$:
over all the networks. We optimize the neural pathway network $\hat{f}_{k}$ for training data points that are closest to prototype $p_{k}$. The training procedure is summarized in Algorithm 1.
Algorithm 1 Neural Pathways Training.
Training data $\mathcal{D}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(x_{j},y_{j})\}_{j=1}^{N}$, no. of prototypes $K\in\mathbb{N}_{+}$, loss function $\ell$.
Discovering Prototypes:
Deepen networks:
    For $k=1,\dots,K$:
     $\hat{f}_{k}\leftarrow\textrm{deepen}(\hat{f}_{k})$
Neural Pathway Training:
   For $k=1,\dots,K$:
     $\hat{f}_{k}\leftarrow\underset{\hat{f}_{k}}{\operatorname*{arg\,min}}\,\underset{\underset{k\in\operatorname*{arg\,min}_{j}\{\|x-p_{j}\|\}}{(x,y)\in\mathcal{D}}}{\sum}\,\ell(\hat{f}_{k}(x),y)$
return Neural pathway parameters $\{\hat{f}_{k}\}_{k=1}^{K}$ and prototype locations $\{p_{k}\}_{k=1}^{K}$.
Inference. At inference time, each test sample $x$ is assigned to its nearest prototype $p_{k}$ where $k\in\operatorname*{arg\,min}_{j\in\{1,\dots,K\}}\{\|x-p_{j}\|\}$ and the prediction is made by the $k$-th neural pathway $\hat{f}_{k}$.
3 Mathematical Analysis of Neural Pathways
We now analyze the approximation capacity of the neural pathways framework, when applied to MLPs with trainable PReLU activation functions.
3.1 Preliminaries
Before presenting our main result, Theorem 3.4, we first formalize the necessary mathematical terminology.

### v2

# Approximation Rates and VC-Dimension Bounds for (P)ReLU MLP Mixture of Experts

@@ Abstract
Mixture-of-Experts (MoEs) can scale up beyond traditional deep learning models by employing a routing strategy in which each input is processed by a single “expert” deep learning model. This strategy allows us to scale up the number of parameters defining the MoE while maintaining sparse activation, i.e., MoEs only load a small number of their total parameters into GPU VRAM for the forward pass depending on the input. In this paper, we provide an approximation and learning-theoretic analysis of mixtures of expert MLPs with (P)ReLU activation functions. We first prove that for every error level $\varepsilon>0$ and every Lipschitz function $f:[0,1]^{n}\to\mathbb{R}$, one can construct a MoMLP model (a Mixture-of-Experts comprising of (P)ReLU MLPs) which uniformly approximates $f$ to $\varepsilon$ accuracy over $[0,1]^{n}$, while only requiring networks of $\mathcal{O}(\varepsilon^{-1})$ parameters to be loaded in memory. Additionally, we show that MoMLPs can generalize since the entire MoMLP model has a (finite) VC dimension of $\tilde{O}(L\max\{nL,JW\})$, if there are $L$ experts and each expert has a depth and width of $J$ and $W$, respectively.
@@ 1 Introduction
With the advent of large foundation models, the need to scale deep learning models beyond what is feasible to process on a single machine has become exceedingly relevant. Mixture of Experts (MoE) models present a solution to this problem via a sparse activation strategy. In MoEs, every input is first routed to one amongst a large number of expert deep learning models and then processed therewith. This allows MoEs to scale up, while maintaining a low/constant computational cost during the forward pass since only a subset of the overall model needs to be loaded into GPU video random-access memory (VRAM) for any given input. This has led to MoEs such as Mixtral (Jiang et al., 2024), Gemini (Google, 2023), and several others, e.g. (Jacobs et al., 1991; Jordan & Xu, 1995; Meila & Jordan, 2000; Shazeer et al., 2017; Guu et al., 2020; Lepikhin et al., 2021; Fedus et al., 2022; Barham et al., 2022; Majid & Tudisco, 2024), to become a viable solution in scaling up large language models (Radford et al., 2018; Brown et al., 2020). However, the analytical and statistical underpinnings of MoEs in deep learning are comparatively less understood in contrast to empirical investigations.
This paper adds to the theoretical understanding of this subject by studying MoEs whose experts are (small) multilayer perceptrons (MLPs) with (P)ReLU activation function (MoMLPs). A key feature of MoEs is that they can maintain a small/fixed computational cost during the forward pass, for any given input $x\in[0,1]^{n}$, even if the overall model complexity may be large. Our main result (Theorem 4.1) analyzes the complexity of MoEs when uniformly approximating an arbitrary Lipschitz (Lebesgue) almost-everywhere continuously differentiable function $f:[0,1]^{n}\to\mathbb{R}^{m}$ by an MoMLP with (P)ReLU activation function to any prespecified error $\varepsilon>0$. We focus on the trade-off between the maximum number of parameters loaded into VRAM by any expert model $\{\hat{f}_{l}\}_{l=1}^{\ell}$ while predicting from any given input, against the total number of experts required to maintain that constant number of activated parameters in the forward pass. Summarized in Table 1, our main result shows that a constant active complexity in the forward pass can be maintained amongst all experts, but at the cost of an exponentially large number of locally-specialized experts $\{\hat{f}\}_{l=1}^{\ell}$ and regions of specialization $\{C_{l}\}_{l=1}^{\ell}$. Our complexity estimates are approximately optimal as they nearly match the Vapnik-Chervonenkis (VC) lower bounds derived in Shen et al. (2021). That is, the uniform approximation of an arbitrary such $f$ on $[0,1]^{n}$, with an an error of $\varepsilon>0$, requires at least $\Omega(\varepsilon^{-n/2})$ total model parameters (Yarotsky, 2018; Kratsios & Papon, 2022; Shen et al., 2021, 2022b). It is here where MoEs have an advantage since not all of these parameters need to be loaded into active memory for any given input; thus MoEs are genuinely sparsely activated.
From the statistical learning perspective, a key property of MoEs (e.g. the top MoMLP model) is that they can maintain a given level of activation in the forward pass while the entire MoE model can maintain a finite VC-dimension (Theorem 4.2). This is key, for instance, in classification applications, as the fundamental theorem of PAC learning (see e.g. (Shalev-Shwartz & Ben-David, 2014, Theorem 6.7) or the results of Blumer et al. (1989); Hanneke (2016); Brukhim et al. (2022)) implies that such a machine learning model generalizes beyond the training data if and only if it has finite VC dimension.
Summary of Contributions
Table 1 summarizes our main contributions, both to the approximation theory and learning theory of MoE models, in the context of the toy mixture of (P)ReLU MLP models. All results illustrate the trade-off between individual (expert) complexity and the complexity shared across the set of experts when uniformly approximating a target $\alpha$-Hölder function, where $0<\alpha\leq 1$. Note that $\alpha=1$ is equivalent to the target function being differentiable (Lebesgue) almost-everywhere.
Our approximation theoretic result (Theorem 4.1) records the number of parameters required to perform a uniform approximation on a high-dimensional Euclidean space on $[0,1]^{n}$. Our statistical learning theoretic result (Theorem 4.2) yields a bound on the VC dimension of the entire class of MoEs with just enough approximation power to perform this approximation.
Table 1: Parametric Complexity and VC-Dimension of MoMLP model with no. experts-to-expert complexity parameter $r\in\mathbb{R}$; performing an $\varepsilon>0$ approximation of an $\alpha$-Hölder function $f:[0,1]^{n}\to\mathbb{R}$; for $n\in\mathbb{N}$.
When $r\geq 0$ more model complexity is distributed across many “small experts”. When $r<0$ fewer experts define the MoE and, as a result, each expert MLP must depend on more parameters such that the entire MoE obtain an accurate approximation of the target function.
Parameter Estimate
No. Expert Parameters $\mathcal{O}(\max\{1,\varepsilon^{-r}\})$
No. Experts $\mathcal{O}\big{(}\max\{1,\varepsilon^{2r/n-1/\alpha}\}\big{)}$
Routing Complexity $\mathcal{O}\big{(}\max\{1,(\frac{2r}{n}-\frac{1}{\alpha})\log(\varepsilon)\}\big{)}$
VC Dimension MoE $\tilde{\mathcal{O}}\big{(}\max\{1,\varepsilon^{2r/n-1/\alpha}\}\,\max\{\varepsilon^{2r/n-1/\alpha},\varepsilon^{-r}\}\big{)}$
Observe that, setting $r=\frac{2}{n}$ in Table 1, yields for ReLU MLPs derived in Yarotsky (2017); the optimality of which is expressed in terms of VC dimension in Shen et al. (2022b). Likewise, the VC dimension of the MoMLP is roughly equal to that of ReLU MLPs computed in Bartlett et al. (2019).
2 Related Work
@@
Deep Learning Models with Few Parameters in Active Memory. Deep learning models with highly oscillatory “super-expressive” activation functions (Yarotsky & Zhevnerchuk, 2020; Yarotsky, 2021; Zhang et al., 2022) are known to achieve dimension-free approximation rates, thus effectively require a (relatively) feasible number of parameters to be loaded into VRAM during the forward pass. As we will see in Proposition 4.4, many of these models have an infinite VC dimension even when they are restricted to having a bounded depth and width; see Jiao et al. (2023, Lemma 3.1) for $\operatorname{ReLU-Sin-2^{x}}$-networks. Their unbounded VC dimension implies that the classifiers implemented by these models do not generalize on classification problems. Thus, the real performance of these models does not need to achieve the approximation-theoretic optima since they can only learn from a finite number of noisy training instances. Alternatively, a feasible number of parameters in deep learning models with standard activation functions may be guaranteed by restricting classes of well-behaved target functions such as Barron functions (Barron, 1993), functions of mixed-smoothness (Suzuki, 2019), highly smooth functions (Mhaskar, 1996; Galimberti et al., 2022; Gonon et al., 2023; Opschoor et al., 2022), convex functions (Bach, 2017), functions with compositional structure (Mhaskar et al., 2017), or other restricted classes. However, there are generally no verifiable guarantees that a target function encountered in practice has the necessary structure for these desired approximation theorems to hold.
Universal Approximation in Deep Learning. Several results have recently considered the expression power of deep learning models. These include universal approximation guarantees for MLPs (Cybenko, 1989; Hornik et al., 1989; Lu et al., 2017; Suzuki, 2019; Yarotsky, 2017, 2018; Voigtlaender & Petersen, 2019; Bolcskei et al., 2019; Gühring et al., 2020; De Ryck et al., 2021; DeVore et al., 2021; Daubechies et al., 2022; Kratsios & Zamanlooy, 2022; Zhang et al., 2022; Opschoor et al., 2022; Zamanlooy & Kratsios, 2022; Shen et al., 2022b; Cuchiero et al., 2023; Voigtlaender, 2023; Benth et al., 2023; Mao & Zhou, 2023; Yang & Zhou, 2024), CNNs (Petersen & Voigtlaender, 2020; Yarotsky, 2022), spiking neural networks (Neuman & Petersen, 2024), residual neural networks (Tabuada & Gharesifard, 2021), transformers (Yun et al., 2019, 2020; Kratsios & Papon, 2022; Fang et al., 2023), random neural networks (Gonon et al., 2023), recurrent neural network models (Grigoryeva & Ortega, 2018; Gonon & Ortega, 2021; Hutter et al., 2022; Galimberti et al., 2022; hoon Song et al., 2023), and several others. In each these cases, one typically considers the expressivity of a single “expert” model and not a mixture thereof. Our analysis can be customized to any of these settings to yield analogues of Theorem 4.1.
Foundations of MoEs. MoE models have been heavily studied since their inception. Most results have focused on identifying the correct expert to best route any given input to (Teicher, 1960, 1963; Wang et al., 1996), the construction of effective routing mechanisms (Wang et al., 2017) selection (Wang et al., 1996), MoE training (Larochelle et al., 2009; Akbari et al., 2024), statistical convergence guarantees for classes of MoEs (Chen, 1995; Ho et al., 2022), robustness guarantees for such models (Puigcerver et al., 2022), amongst several other types of guarantees. However, to our knowledge, there are no available approximation guarantees for MoE or VC-dimension bounds for deep-learning-based MoEs. Thus, our results would be adding to the approximation theoretic foundations of MoE models as well as to the statistical foundations of deep-learning-based MoEs.
Prototypes and Partitioning. Each region in our learned partition of the input space is associated with a representative point therein called a prototype. Prototypes (also called landmarks) are routinely used in image classification (Mensink et al., 2012), few-shot learning (Snell et al., 2017; Cao et al., 2020), dimensionality reduction (Law et al., 2019), in complex networks (Keller-Ressel & Nargang, 2023), and geometric deep learning (Ghadimi Atigh et al., 2021) to tractably encode massive structures. They are also standard in classical clustering algorithms such as $K$-medoids or $K$-means, wherein the part associated with each medoid (resp. centroid) defines a Voronoi cell or Voronoi region (Voronoi, 1908). Moreover, while partitioning is commonly employed in deep learning for various purposes, such as proving universal approximation theorems (Yarotsky, 2017; Lu et al., 2021b; Gühring & Raslan, 2021) or facilitating clustering-based learning (Zamanlooy & Kratsios, 2022; Trask et al., 2022; Ali & Nouy, 2023; Srivastava et al., 2022), existing approaches typically involve loading the entire model into VRAM. Our approach, however, differs by relying on a learned partition of the input space, where each part is associated with a distinct small neural network. Importantly, the complete set of networks forming the MoMLPs does not need to be simultaneously loaded into VRAM during training or inference.
Paper Overview.
Our paper is organized as follows. Section 3 contains preliminary notation, definitions, and mathematical background required for the formulation of our main results. Section 4 contains our main approximation (Theorem 4.1) and learning theoretic (Theorem 4.2) guarantees. Section 5 dives into the details of why MoEs can achieve arbitrary precision while maintaining a feasible active computational complexity by explaining the derivation of our main approximation theorem; the details of which are relegated to Appendix B. A technical version (Theorem 5.3) of our main approximation guarantee is then presented, which allows for the approximation of continuous functions of arbitrarily low regularity and for the organization of the experts defining the MoMLP into more general tree structures. Appendix D contains technical derivations of our main results as well as experimental elucidation of the benefit of MoEs, and specifically the toy MoMLP model.
3 Preliminaries
We standardize our notation, define the necessary mathematical formalisms to state our main results and define our toy MoE Model.
3.1 Notation.